OpenStudy (anonymous):

Which of the following is equal to (2x/3 - 7) + 7 A. (2x - 7) + 21 B. 2x - 21/3 C. 3x/2 D. 2x/3 @texaschic101

hero (hero):

\(\left(\frac{2x}{3} - 7\right) + 7 = \frac{2x}{3} + \left(-7 + 7\right)\)

hero (hero):

@sorour you can't multiply both sides by 3 if there's no equal sign.

hero (hero):

It's only one expression.

OpenStudy (anonymous):

I need an explanation I do not know how to do this I suck at fractions

hero (hero):

You have to shift the parentheses.

OpenStudy (anonymous):

I'm confused

hero (hero):

Associative property: (a + b) + c = a + (b + c)

OpenStudy (anonymous):

@Hero yeah ur right

OpenStudy (texaschic101):

I missed that....I was trying to multiply by 3 as well....good one Hero

OpenStudy (texaschic101):

(2x/3 (-7 + 7) gives you what sirryan ? Is that right Hero ?

hero (hero):

You have to include the plus between otherwise you make it look like distributive property

hero (hero):

2x/3 + (-7 + 7)

OpenStudy (texaschic101):

gotcha :)

OpenStudy (texaschic101):

sirryan, what happens to the 7's ?

hero (hero):

It's definitely a tricky one.

OpenStudy (anonymous):

They crossout

OpenStudy (texaschic101):

yes, but you were smart enough to figure it out

OpenStudy (texaschic101):

they do cross out, leaving you with what ?

OpenStudy (anonymous):

2x/3

OpenStudy (texaschic101):

yes

OpenStudy (texaschic101):

medal goes to Hero :)

OpenStudy (anonymous):

So I'm confused on my answer

OpenStudy (texaschic101):

answer D

OpenStudy (texaschic101):

your 7's cancel out leaving you with just 2x/3

hero (hero):

It is more clearer to see if you did this: (2x/3 - 7) + 7 = (-7 + 2x/3) + 7 = 7 + (-7 + 2x/3) Then you would be able to simply shift parentheses = (7 - 7) + 2x/3 = 0 + 2x/3 = 2x/3